# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ 3.3 Discussion
We also qualitatively evaluated our framework in an informal study with two US board-certified dermatologists familiar with the labeling tool (Liu et al., 2020). Specifically, we discussed individual cases with particularly low annotation certainty by showing input images alongside meta information (sex, age, etc.) and the corresponding annotations (cf. Figure 5). Discussing these cases takes considerable time while the dermatologists try to understand how the annotators came to their respective conclusions. In most cases, the disagreement was attributed to inherent uncertainty, i.e., missing information, inconclusive images, etc. In only few cases, the disagreement was attributed to annotator mistakes or annotation quality in general – e.g., inexperienced annotators, annotators ignoring meta information etc. Again, this highlights the difficulty of disentangling annotation and inherent uncertainty in cases with high disagreement (as discussed in Sections 2.2 and 2.3) and is in line with related work on “meta-annotation” for understanding sources of disagreement (Sandri et al., 2023; Bhattacharya et al., 2019). However, this also highlights that our uncertainty-adjusted metrics appropriately take both sources of uncertainty into account.
@@
The results for this case study indicate, using our annotation certainty measure, that a large portion of the dataset exhibits high ground truth uncertainty. The current approach (Liu et al., 2020) of deterministically aggregating annotations using IRN and then evaluating against the corresponding top-1 labels largely ignores this uncertainty. In our framework, using the PrIRN model, this implicitly corresponds to evaluation at infinite reliability, i.e., full trust in all annotators. Instead, our approach to evaluation paints a more complete picture by computing uncertainty-adjusted (top-k) accuracy across a range of reliabilities, corresponding to different trust “scenarios”. In practice, this not only allows to compare models across these different scenarios but also highlights the expected variation in performance. Moreover, we show that performance is always relative to the chosen aggregation model, as highlighted using our alternative Plackett-Luce (PL) (Plackett, 1975; Luce, 2012) based model. As feedback from dermatologists indicates that most of this ground truth uncertainty stems from inherently ambiguous cases, we also explored metrics considering more than the top-1 conditions for evaluation. Here, performance drops rather drastically, indicating potential negative consequences for patients when lower ranked conditions are not appropriate. For example, seemingly random conditions on the 2nd or 3rd place of the prediction set can easily lead to confusion or anxiety. Overall, we believe that our framework will help with model development and make model selection more robust and thereby positively influence patient outcome.
@@ 4 Related work
Annotator disagreement has been discussed extensively and early on in medicine (Feinstein and Cicchetti, 1990; McHugh, 2012; Raghu et al., 2019; Schaekermann, 2020) as well as machine learning (Dawid and Skene, 1979; Smyth et al., 1994). Natural language processing, for example, has particularly strong work on dealing with disagreement (Reidsma and op den Akker, 2008; Aroyo and Welty, 2014, 2015; Schaekermann et al., 2016; Dumitrache et al., 2019; Röttger et al., 2022; Abercrombie et al., 2023), see (Pavlick and Kwiatkowski, 2019) for an overview. As crowdsourcing human annotations has become a standard tool in creating benchmarks across the field (Kovashka et al., 2016; Sorokin and Forsyth, 2008; Snow et al., 2008) – though not without criticism (Röttger et al., 2021) – most work focuses on resolving or measuring disagreement and aggregating annotations. Methods for measuring disagreement (Feinstein and Cicchetti, 1990; Powers, 2012; Uma et al., 2021) are often similar across domains. However, measures such as Fleiss’/Cohen’s kappa (Cohen, 1960; Fleiss et al., 2003), percent agreement (McHugh, 2012), or intra-class correlation coefficient (Landis and Koch, 1977) are only applicable to annotations with single class responses such that generalized approaches (Braylan et al., 2022) or custom measures are used for more structured annotations (Pavlick and Kwiatkowski, 2019). Resolving disagreement is typically done computationally (e.g., through majority vote). However, recent work has explored domain-specific interactive approaches for resolving disagreement, involving discussions or deliberation (Schaekermann et al., 2020b, a, 2019a; Schaekermann, 2020; Pradhan et al., 2022; Drapeau et al., 2016; Chen et al., 2019b; Silver et al., 2021; Bakker et al., 2022) or relabeling (Sheng et al., 2008), and reducing disagreement by co-designing labeling with experts (Freeman et al., 2021). Recent work also considers properly modeling disagreement (Vitsakis et al., 2023) or performing meta analysis (Sandri et al., 2023; Bhattacharya et al., 2019), trying to understand sources of disagreement. For benchmarks, disagreement is generally addressed by aggregating labels from multiple annotators to arrive at what is assumed to be the single correct label. This can involve basic majority voting or more advanced methods (Dawid and Skene, 1979; de Marneffe et al., 2012; Pham et al., 2017; Warby et al., 2014; Carvalho and Larson, 2013; Gaunt et al., 2016; Tian et al., 2019), including inverse rank normalization (IRN) as discussed in this paper. Often, aggregation is also performed using probabilistic models from the crowdsourcing and truth-discovery literature (Yin et al., 2008; Welinder and Perona, 2010; Li et al., 2012; Dong et al., 2009; Zhao et al., 2012; Wang et al., 2012; Bachrach et al., 2012; Rodrigues and Pereira, 2018; Guan et al., 2018; Chu et al., 2021; Gordon et al., 2022), see (Yan et al., 2014; Zheng et al., 2017) for surveys. However, evaluation is often based on point estimates and the impact of annotator disagreement on evaluation is generally poorly understood (Gordon et al., 2021). While, e.g., (Reidsma and op den Akker, 2008; Collins et al., 2022) train models on individual annotators, (de Marneffe et al., 2012; Nie et al., 2020) perform evaluation on aggregated probabilities instead of top-1 labels, and (Gao et al., 2017) trains on label distributions rather than discrete aggregated labels, there is no common understanding for dealing with annotation uncertainty for evaluation. Instead, following (Plank, 2022), disagreement is often treated as label noise. Here, early work (Angluin and Laird, 1987; Kearns, 1998; Kearns and Li, 1993; Lawrence and Schölkopf, 2001) assumes uniform or class-conditional label noise, while more recent work (Beigman and Klebanov, 2009; Oyen et al., 2022) also considers feature-dependent or annotator-dependent noise. Popular methods try to estimate the label noise distributions (Northcutt et al., 2019; Hendrycks et al., 2018) in order to prune or re-weight examples. Such approaches have also been utilized to infer annotator confusion matrices (Zhang et al., 2020; Tanno et al., 2019), similar to annotator quality in crowdsourcing. We refer to (Chen et al., 2019a; Zhang et al., 2022) for good overviews and note that there is also some similarity to partial label learning (Hüllermeier and Beringer, 2005; Nguyen and Caruana, 2008; Cour et al., 2011; Wang et al., 2022).
@@
Overall, work on handling annotation disagreement is very fragmented, as highlighted above and in recent surveys (Uma et al., 2021). Moreover, many works treat symptoms such as label noise rather than treating annotator disagreement as uncertainty in the ground truth. This is emphasized in recent position papers (Baan et al., 2022; Plank, 2022; Basile et al., 2021) that argue for common frameworks to deal with this challenge. Indeed, recent work (Sculley, 2007; Maier-Hein et al., 2018; Cabitza et al., 2020; Belz et al., 2023) demonstrates that many results in machine learning are not reproducible, in part due to annotation uncertainty. This has also been the basis for several workshops and challenges on directly learning with disagreement (Leonardelli et al., 2023). Closest to our work, (Gordon et al., 2021; Lovchinsky et al., 2020) propose methods to incorporate label disagreement into evaluation metrics. However, their work is limited to binary classification tasks. Moreover, the considered annotations are unstructured, i.e., annotators merely provide single labels. In contrast, our framework for evaluation with annotation uncertainty if independent of task, domain or annotation format. In contrast to our findings, (Chen et al., 2021) argues that label noise in validation sets can still lead to reliable model selection on rather unambiguous datasets such as CIFAR10. Finally, (Collins et al., 2022) evaluates against labels aggregated from random subsets of annotators on CIFAR10-H, which can be seen as a bootstrapping approach to statistical aggregation in our framework.
@@ 5 Conclusion
In this paper, we proposed a framework for evaluating AI models under uncertain ground truth. We believe that ground truth uncertainty stems from annotation uncertainty as well as inherent uncertainty and is typically observed in terms of annotator disagreement: in almost all supervised learning tasks, ground truth labels are implicitly or explicitly obtained by aggregating annotations, e.g., counting frequencies and majority voting. Unfortunately, this type of deterministic aggregation typically ignores the underlying uncertainty which can have severe consequences for safety-critical applications such as health. Instead, we introduce a framework based on a statistical model for aggregating annotations that explicitly accounts for uncertainty. Further, we propose a novel measure of annotation uncertainty and present uncertainty-adjusted metrics for evaluating and comparing AI systems. Applied to a case study in skin condition classification, our framework allowed us to make several important observations that previous work (Liu et al., 2020) missed: First, a large portion of cases exhibits high ground truth uncertainty which, according to dermatologists, often stems from inherent uncertainty. Second, classifier performance often degrades and exhibits significant variation under our uncertainty-adjusted metrics. Third, classifiers perform poorly when taking into account inherent uncertainty by evaluating not only against possible top-1 ground truth labels. Our framework can readily be applied to other settings by adapting the statistical aggregation model to the annotations at hand. We believe that properly accounting for ground truth uncertainty in evaluation will play a critical role in successfully tackling more nuanced and ambiguous tasks.
@@ Acknowledgements
We would like to thank Annisah Um’rani and Peggy Bui for their support of this project as well as Naama Hammel, Boris Babenko, Katherine Heller, Verena Rieser and Dilan Gorur for their feedback on the manuscript.
@@ Appendix A Further introductory examples and details
Toy dataset and CIFAR10: For our toy dataset as well as CIFAR10, we model $p(\lambda|b,x)=\mathcal{D}(\lambda;\mu)$ as Dirichlet distribution with concentration parameter vector $\mu$. We define $\mu=\gamma s+\alpha\mathbf{1}$ where $s$ is a vector of sufficient statistics with the $c$’th element $s_{c}$ denoting the number of annotators that have chosen $c$ as (top-1) label, $\gamma>0$ is the reliability parameter, $\alpha>0$ is a small prior parameter that we assign to each label $c$, and $\mathbf{1}$ a vector of ones of the same shape as $s$. As shown in Figure 2, a higher reliability parameter $\gamma$ makes the plausibilities $\lambda\sim p(\lambda|b)$ more concentrated, resulting in higher annotation certainty. Here, the effect of $\alpha$ vanishes with higher reliability or more annotators; essentially, $\alpha$ represents a minimum number of annotators.
@@
Figure 9: Annotation certainty for 1k examples from the Wikipedia toxicity dataset (Wulczyn et al., 2017) as well as those examples identified by the majority of annotators as toxic or very toxic. On the right, we show an example score distribution and samples from the corresponding posterior over plausibilities. Certainty is low for approximately $50\%$ of the toxic examples.
@@
Wikipedia Toxicity (Wulczyn et al., 2017) is a natural language dataset for detecting toxicity in comments posted in Wikipedia page discussions. Around 10 annotators rate each comment between -2 (very toxic) and 2 (very healthy). Then, a comment can be treated as toxic if the mean rating is below 0. This suggests that the scores are assumed to be normally distributed. In this case, we define the plausibilities as the mean and standard deviation: $\lambda=(\mu,\sigma^{2})$. As before, plausibilities parameterize the distribution over targets. Choosing an appropriate prior [^3] We can put a normal-inverse-gamma prior on $\mu$ and $\sigma^{2}$, i.e., $\mu\sim\mathcal{N}(\mu_{0},\nu)$ and $\sigma^{2}\sim\Gamma(\alpha,\beta)$ and estimate the parameters $\mu_{0},\nu,\alpha$ and $\beta$ using the scores. for $\mu$ and $\sigma$ allows to calculate certainty as the maximum of the two possible outcomes: $\mathcal{N}(y\leq 0;\mu,\sigma^{2})$ and $\mathcal{N}(y>0;\mu,\sigma^{2})$. Figure 9 shows the computed certainties for the first 1k examples of the dataset. A significant part of the examples is highly uncertain, especially among the toxic cases (majority of annotators responding -2 or -1, roughly 1% of all cases). We also show one qualitative example with the corresponding annotations and plausibility samples. Note that, in this case, plausibilities correspond to one-dimensional Gaussian distributions which are plotted using mean and precision.
@@
AVA (Murray et al., 2012) is a dataset of visual aesthetics scores. Photography enthusiasts and experts rated more than 200k images on a scale between 1 and 10. Compared to CIFAR10, this is a more subjective and thus ambiguous labeling task. Nevertheless, the dataset is commonly used for predicting visual aesthetics, e.g., by predicting poor or good images (Lu et al., 2014; Esfandarani and Milanfar, 2018). For simplicity, this can be done by taking the mean of the ratings and classifying images with mean score lower than $5$ as poor and those with mean score higher or equal to $5$ as good. This suggests that the scores are assumed to be normally distributed. Specifically, $\lambda=(\mu,\sigma^{2})$, modeling mean and standard deviation. Using the same normal-inverse-gamma prior as above, we can compute certainty by estimating $\mathcal{N}(y\leq 5;\mu,\sigma^{2})$ and $\mathcal{N}(y>5;\mu,\sigma^{2})$. Figure 10 shows the computed certainties as well as an example of a typical score distribution and plausibility samples. Surprisingly, for the first 10k images of AVA, most examples receive high certainty. However, this does not imply that there is no inherent ambiguity, as highlighted in Figure 10. Instead, it might just show that there is lower annotation uncertainty due to the high number of, on average, $222$ annotations.
@@
Figure 10: Left: Annotation certainty for the first 10k AVA (Murray et al., 2012) images considering a binary classification setting of poor images (score $<5$) against good images (score $\geq 5$). Certainty might be high primarily due to the large number of annotators. Right: Score distribution for one example together with samples from the corresponding posterior $p(\lambda=(\mu,\sigma^{2})|b)$.
Appendix B Methods for differential diagnoses in health
Having presented our general framework in Section 2, we present in this section technical details and specific methods for applying our methodology to differential diagnosis as observed in our dermatology case study. As indicated in Section 3, we model differential diagnoses as partial rankings and discuss meaningful aggregation models $p(\lambda|b,x)$ that facilitate sampling plausibilities. Specifically, we follow (Liu et al., 2020), and use inverse rank normalization (IRN) as deterministic aggregation as well as propose a probabilistic variant. We also show how a Plackett–Luce model (Plackett, 1975; Luce, 2012) can be adapted to this case. Remember that in this paper we assume $p(\lambda|b,x)=p(\lambda|b)$ in our aggregation schemes.
B.1 Modeling annotations as partial rankings
We model differential diagnoses as partial rankings. In our case study, for each case, each annotator selects a subset of conditions they think might be the true diagnosis. In addition, they assign a confidence score to each condition, indicating how likely they think that condition is the true diagnosis. The subsets of conditions selected by different annotators may differ in size. Moreover, the confidence scores are usually not comparable across annotators which is why they are instead used to produce a ranking of the conditions for each annotator (Liu et al., 2020). This results in partial rankings because there may be ties among conditions (i.e., conditions with the same confidence score) and the large majority of possible conditions remains unranked. Formally, we write
$b=(b_{{}{}{1}}\succ b_{{}{}{2}}\ldots\succ b_{{}{}{L}}).$ (13)
We call groups of tied conditions “blocks”, and each block contains one or more labels, $b_{{}{}{i}}=\{{\sigma_{{i}{1}}},\ldots,{\sigma_{{i}{|b_{{}{}{i}}|}}}\}$ with ${\sigma_{{i}{j}}}\in[K]$ and $K$ being the number of classes. Here, $b_{{}{}{i}}\succ b_{{}{}{j}}$ indicates that classes in the $i$-th block are ranked higher than those in the $j$-th block (according to the confidence scores). We assume the blocks are non-empty, mutually exclusive, i.e., $\emptyset\neq b_{{}{}{i}}\subseteq[K]$ and $b_{{}{}{i}}\cap b_{{}{}{j}}=\emptyset$, and that $\bigcup_{i=1}^{L}b_{i}=[K]$. Note that there might exist multiple equivalent permutations within the individual blocks $b_{i}$ such that we use $\mathcal{S}(b)$ to denote the set of all permutations compatible with Equation (13). In the literature discussing ranking metrics and aggregation (Sakai, 2013; Webber et al., 2010; Sculley, 2007; Wu and Crestani, 2003), partial rankings are also commonly referred to as top-$k$ rankings with ties. While some metrics such as Kendall’s tau and Spearman’s footrule have been adapted to ties (Fagin et al., 2004; Vigna, 2015), modeling partial rankings is generally non-trivial. This also holds for any aggregation model $p(\lambda|b,x)$ as we will see in Section B.4.
[…]
We interpret the above example as follows: the annotator declares that Hemangioma is their first choice, followed by either Melanocytic Nevus, Melanoma, or O/E. However, for the latter three conditions, the annotator is unable to say which one is more likely than the others. That is, these three conditions are tied for the second rank. Implicitly, this means that all $3!=6$ possible permutations would be equivalent:
[…]
$\ldots$
[…]
The larger the individual blocks, the larger the number of compatible permutations becomes. Also, annotators usually ignore a large set of conditions. Strictly speaking, the ignored conditions can be thought of as another block, such that
[…]

### v2

@@ 3.4 Discussion
The results in this paper indicate, using our annotation certainty measure, that a large portion of the dataset exhibits high ground truth uncertainty. The current approach [58] of deterministically aggregating annotations using IRN and then evaluating against the corresponding top-1 labels largely ignores this uncertainty. In our framework, using the PrIRN model, this implicitly corresponds to evaluation at infinite reliability, i.e., full trust in all annotators. Instead, our approach to evaluation paints a more complete picture by computing uncertainty-adjusted (top-k) accuracy across a range of reliabilities, corresponding to different trust “scenarios”. In practice, this not only allows to compare models across these different scenarios but also highlights the expected variation in performance. Moreover, we show that performance is always relative to the chosen aggregation model, as highlighted using our alternative PL-based models. Differences between results obtained using PrIRN and PL stem from different assumptions or biases. For example, PrIRN’s treatment of ties appears unintuitive, suggesting PL is more reliable for our application. This also leads to generally lower performance when evaluating against PrIRN compared to PL as plausibility mass is more equally distributed for many ambiguous examples due to these ties.
Another important aspect of our uncertainty-adjusted metrics is that no model may be able to achieve, e.g., 100% top-1 uncertainty-adjusted accuracy if the ground truth is deemed sufficiently uncertain. As our uncertainty-adjusted metrics are comparable to their regular counterparts, this indicates that the top-1 prediction is not sufficient. Analogously, we also show that considering more than the top-1 condition for evaluation can be important. Our results show performance drops rather drastically, indicating potential negative consequences for patients when critical lower-ranked conditions are not considered. For example, seemingly random conditions on the 2nd or 3rd place of the prediction set can lead to confusion or anxiety.
We also qualitatively evaluated our framework in an informal study with two US board-certified dermatologists familiar with the labeling tool [58]. Specifically, we discussed individual cases with particularly low annotation certainty by showing input images alongside meta information (sex, age, etc.) and the corresponding annotations (cf. Figure 5). Discussing these cases takes considerable time as the dermatologists try to understand how the annotators came to their respective conclusions. In most cases, the disagreement was attributed to inherent uncertainty, i.e., missing information, inconclusive images, etc. In only a few cases, the disagreement was attributed to annotator mistakes or annotation quality in general – e.g., inexperienced annotators, annotators ignoring meta information etc. Once again, this highlights the difficulty of disentangling annotation and inherent uncertainty in cases with high disagreement (as discussed in Sections 2.2 and 2.4), aligning with related work on “meta-annotation” for understanding sources of disagreement [87, 12]. However, it also emphasizes that our uncertainty-adjusted metrics effectively consider both sources of uncertainty.
Overall, we believe that our framework will help model development and improve model selection robustness, potentially benefiting patient outcomes. We also emphasize that annotation certainty and uncertainty-adjusted metrics are not dependent on the exact task and aggregation model used. Thus, these can readily be applied to other tasks with different annotation formats. Additionally, although PL was motivated here by our dermatology application, it is applicable across a wide range of diagnosis tasks where annotations often come in the form of partial rankings (i.e., differential diagnoses).
@@ 4 Related work
Annotator disagreement has been discussed extensively and early on in medicine [34, 63, 79, 88] as well as machine learning [24, 98]. Natural language processing, for example, has particularly strong work on dealing with disagreement [80, 3, 4, 89, 29, 83, 1], see [71] for an overview. As crowdsourcing human annotations has become a standard tool in creating benchmarks across the field [51, 100, 99] – though not without criticism [82] – most work focuses on resolving or measuring disagreement and aggregating annotations. Methods for measuring disagreement [34, 77, 104] are often similar across domains. However, measures such as Fleiss’/Cohen’s kappa [20, 36], percent agreement [63], or intra-class correlation coefficient [54] are only applicable to annotations with single class responses such that generalized approaches [13] or custom measures are used for more structured annotations [71]. Resolving disagreement is typically done computationally (e.g., through majority vote). However, recent work has explored domain-specific interactive approaches for resolving disagreement, involving discussions or deliberation [93, 92, 90, 88, 78, 27, 18, 97, 8] or relabeling [95], and reducing disagreement by co-designing labeling with experts [38]. Recent work also considers properly modeling disagreement [107] or performing meta analysis [87, 12], trying to understand sources of disagreement.
@@
For benchmarks, disagreement is generally addressed by aggregating labels from multiple annotators to arrive at what is assumed to be the single correct label. This can involve basic majority voting or more advanced methods [24, 25, 73, 110, 16, 40, 103], including inverse rank normalization (IRN) as discussed in this paper. Often, aggregation is also performed using probabilistic models from the crowdsourcing and truth-discovery literature [117, 112, 57, 26, 120, 108, 7, 81, 43, 19, 42], see [116, 121] for surveys. However, evaluation is often based on point estimates and the impact of annotator disagreement on evaluation is generally poorly understood [41]. While, e.g., [80, 21] train models on individual annotators, [25, 67] perform evaluation on aggregated probabilities instead of top-1 labels, and [39] trains on label distributions rather than discrete aggregated labels, there is no common understanding for dealing with annotation uncertainty for evaluation. Instead, following [76], disagreement is often treated as label noise. Here, early work [2, 48, 49, 55] assumes uniform or class-conditional label noise, while more recent work [10, 70] also considers feature-dependent or annotator-dependent noise. Popular methods try to estimate the label noise distributions [68, 44] in order to prune or re-weight examples. Such approaches have also been utilized to infer annotator confusion matrices [119, 102], similar to annotator quality in crowdsourcing. We refer to [17, 118] for good overviews and note that there is also some similarity to partial label learning [45, 66, 22, 109].
@@
Overall, work on handling annotation disagreement is very fragmented [104] and focused on treating symptoms such as label noise rather than the underlying uncertainty in the ground truth. This is emphasized in recent position papers [6, 76, 9] that argue for common frameworks to deal with this challenge. Indeed, recent work [94, 62, 14, 11] demonstrates that many results in machine learning are not reproducible, in part due to annotation uncertainty. This has also been the basis for several workshops and challenges [56] and the main motivation for this paper. Closest to our work, [41, 59] propose methods to incorporate label disagreement into evaluation metrics. However, their work is limited to binary classification tasks. Moreover, the considered annotations are unstructured, i.e., annotators merely provide single labels. In contrast, our framework for evaluation with annotation uncertainty if independent of task, domain or annotation format. Finally, [21] evaluates against labels aggregated from random subsets of annotators, which can be seen as a bootstrapping approach to statistical aggregation in our framework.
@@ 5 Conclusion
In almost all supervised learning tasks, ground truth labels are implicitly or explicitly obtained by aggregating annotations, e.g., majority voting. These deterministic aggregation methods ignore the underlying uncertainty and tend to over-estimate performance of models. This can be particularly critical when assessing the quality of health models.
In this paper, we propose a statistical model for aggregating annotations that explicitly accounts for ground truth uncertainty. Based on this model, we provide a novel framework for understanding and quantifying annotation ambiguity. Specifically, we introduce a measure of annotation certainty to identify ambiguous data points and develop uncertainty-adjusted performance metrics for more reliable AI system evaluation.
We have demonstrated our methodology on a challenging skin condition classification problem addressed in [58]. In this context, our annotation uncertainty measure has revealed that a large proportion of cases in the training and test sets exhibit high ground truth uncertainty which has been ignored previously. Moreover, we observe that classifier performance, as estimated using our uncertainty-adjusted metrics, can exhibit significant variations, hinting at the possibility that true performance may be much lower in a real deployment scenario. The variations in model performance estimates become more pronounced when evaluating not only against possible top-1 ground truth labels but prediction sets derived from differential diagnoses (as highlighted using set accuracy).
While we have focused on a dermatology application, the framework’s versatility allows adaptation to other research domains, with direct relevance to diagnosis tasks involving differential diagnoses or partial rankings. This approach is especially valuable in medical and healthcare contexts where nuanced interpretation of results is crucial.
@@ Acknowledgements
We would like to thank Annisah Um’rani and Peggy Bui for their support of this project as well as Naama Hammel, Boris Babenko, Katherine Heller, Verena Rieser, and Dilan Gorur for their feedback on the manuscript.
@@ Appendix A Further introductory examples and details
Fig. 9: We plot annotation certainty and uncertainty-adjusted (UA) accuracy for CIFAR10-H [52, 72], highlighting the 1000 examples with lowest annotation certainty. There is a small set of $178$ images with certainty below $99\%$ for which annotators tend to disagree. Given the large number of annotators ($50$), this likely indicates examples with inherent uncertainty, as shown for two examples on the right (images and histogram of annotations).
@@
CIFAR10: To move from the toy example in Section 2.1 to a real benchmark, we consider CIFAR10 [52] with the human annotations from [72]. Again, each annotator provides a single label such that we can apply the same methodology. Specifically, we model $p(\lambda|b,x)=\mathcal{D}(\lambda;\mu)$ as Dirichlet distribution with concentration parameter vector $\mu$. We define $\mu=\gamma s+\alpha\mathbf{1}$ where $s$ is a vector of sufficient statistics with the $c$’th element $s_{c}$ denoting the number of annotators that have chosen $c$ as (top-1) label, $\gamma>0$ is the reliability parameter, $\alpha>0$ is a small prior parameter that we assign to each label $c$, and $\mathbf{1}$ a vector of ones of the same shape as $s$. A higher reliability parameter $\gamma$ makes the plausibilities $\lambda\sim p(\lambda|b)$ more concentrated, resulting in higher annotation certainty. Here, the effect of $\alpha$ vanishes with higher reliability or more annotators; essentially, $\alpha$ represents a minimum number of annotators.
Figure 9 (left) reports the corresponding annotation certainty alongside an uncertainty-adjusted version of accuracy. As can be seen, annotators agree for the majority of examples, resulting in certainties close to $1$. Nevertheless, on roughly $178$ test examples, certainty is below $0.99$. As there are 50 annotators per example with only 10 classes, it is likely that these are actually difficult cases, i.e. cases with inherent uncertainty, as confirmed in Figure 9. This corresponds to roughly $0.2\%$ of the examples. Interestingly, recent improvements in accuracy on CIFAR10 are often smaller than $0.2\%$ [^9] According to Papers with Code, https://paperswithcode.com/sota/image-classification-on-cifar-10.. To measure accuracy while taking annotation uncertainty into account, we measure how often the original CIFAR10 ground truth labels coincide with the top-1 labels obtained from sampled plausibilities. We call this uncertainty-adjusted accuracy (formally defined in Section 2.5) and, unsurprisingly, observe that the original labels of [52] (when interpreted as predictions) often perform poorly for examples with high annotation uncertainty (red in Figure 9).
@@
Fig. 10: Annotation certainty for 1k examples from the Wikipedia toxicity dataset [115] as well as those examples identified by the majority of annotators as toxic or very toxic. On the right, we show an example score distribution and samples from the corresponding posterior over plausibilities. Certainty is low for approximately $50\%$ of the toxic examples.
@@
Wikipedia Toxicity [115] is a natural language dataset for detecting toxicity in comments posted in Wikipedia page discussions. Around 10 annotators rate each comment between -2 (very toxic) and 2 (very healthy). Then, a comment can be treated as toxic if the mean rating is below 0. This suggests that the scores are assumed to be normally distributed. In this case, we define the plausibilities as the mean and standard deviation: $\lambda=(\mu,\sigma^{2})$. As before, plausibilities parameterize the distribution over targets. Choosing an appropriate prior [^10] We can put a normal-inverse-gamma prior on $\mu$ and $\sigma^{2}$, i.e., $\mu\sim\mathcal{N}(\mu_{0},\nu)$ and $\sigma^{2}\sim\Gamma(\alpha,\beta)$ and estimate the parameters $\mu_{0},\nu,\alpha$ and $\beta$ using the scores. for $\mu$ and $\sigma$ allows to calculate certainty as the maximum of the two possible outcomes: $\mathcal{N}(y\leq 0;\mu,\sigma^{2})$ and $\mathcal{N}(y>0;\mu,\sigma^{2})$. Figure 10 shows the computed certainties for the first 1k examples of the dataset. A significant part of the examples is highly uncertain, especially among the toxic cases (majority of annotators responding -2 or -1, roughly 1% of all cases). We also show one qualitative example with the corresponding annotations and plausibility samples. Note that, in this case, plausibilities correspond to one-dimensional Gaussian distributions which are plotted using mean and precision.
@@
AVA [65] is a dataset of visual aesthetics scores. Photography enthusiasts and experts rated more than 200k images on a scale between 1 and 10. Compared to CIFAR10, this is a more subjective and thus ambiguous labeling task. Nevertheless, the dataset is commonly used for predicting visual aesthetics, e.g., by predicting poor or good images [60, 31]. For simplicity, this can be done by taking the mean of the ratings and classifying images with mean score lower than $5$ as poor and those with mean score higher or equal to $5$ as good. This suggests that the scores are assumed to be normally distributed. Specifically, $\lambda=(\mu,\sigma^{2})$, modeling mean and standard deviation. Using the same normal-inverse-gamma prior as above, we can compute certainty by estimating $\mathcal{N}(y\leq 5;\mu,\sigma^{2})$ and $\mathcal{N}(y>5;\mu,\sigma^{2})$. Figure 11 shows the computed certainties as well as an example of a typical score distribution and plausibility samples. Surprisingly, for the first 10k images of AVA, most examples receive high certainty. However, this does not imply that there is no inherent ambiguity, as highlighted in Figure 11. Instead, it might just show that there is lower annotation uncertainty due to the high number of, on average, $222$ annotations.
@@
Fig. 11: Left: Annotation certainty for the first 10k AVA [65] images considering a binary classification setting of poor images (score $<5$) against good images (score $\geq 5$). Certainty might be high primarily due to the large number of annotators. Right: Score distribution for one example together with samples from the corresponding posterior $p(\lambda=(\mu,\sigma^{2})|b)$.
Appendix B Plackett–Luce likelihood and Gibbs sampling